what is the formula for a capacity

Grade:12

3 Answers

raj mishra
11 years ago

C=Q/V

C= capacitance

Q= total  Charge on the capacitor

V= Potential difference across Capacitor

shreya shukla
11 years ago

the formula for capacaitance c=q/v can be misleading, as it shows that capacitance changes with q and v. but c is a property of material, just as specific heat is.

 for a parellel plate capacitor, c = εka/d

for spherical capacitor , c=4Πεr, where r is the radius, this suggests that capacitance c depends on dimensional properties only.

akash luckey aggara
11 years ago

c=q/v